OpenStudy (anonymous):

A car drives at a constant speed of 14 m/s around a circle with radius 55 m. What is the centripetal acceleration of the car? A. 2.3 m/s2 B. 0.25 m/s2 C. 3.6 m/s2 D. 3.9 m/s2

OpenStudy (abmon98):

Centripetal acceleration=v^2/r=(14)^2/55=3.56m/s^2